2 Khaleel Compagny produces three products A, B and C. During the year the joint costs of processing the coffee were SAR270,000. Production and sales value information were as follows: Sales Value Product Units at Split-Off Separable Costs Selling Price A 300,000 SAR9 per unit SAR5.00 per unit SAR32 per unit B 200,000 SAR8 per unit SAR3.00 per unit SAR30 per unit C 400,000 SAR7 per unit SAR2.00 per unit SAR20 per unit.

Chose one method to allocate joint costs and allocate the joint costs
Answer #1
Answer
Note- Using the physical output method
JOINT COST = $270000
Total units = 300000 + 200000 + 400000 = 900,000 units
Allocation based on physical production of units =
Product A = $270000 × (300000/9000000) = $90,000
Product B = $270000 × (200000/9000000) = $60,000
Product C = $270000 × (400000/9000000) = $120,000